a cube of side b has a charge q at each of its vertices. Determine the electric field and potential at its centre.

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
electric field at the center will be zero because all components will balance each others.
for electric potential:
body diagonal of the cube =sqrt(3)*b
so distance of each charge from center=half of body diagonal=sqrt(3)*b/2
so potential=8*k*q/r=16*k*q/(sqrt(3)*b)

Sher mohammad
B.tech IIT delhi
faculty Askiitians